At Central Park Zoo, the ratio of sea lions to penguins is 4:11. If there are 84 more penguins than sea lions, how many sea lions are there? Thank you!!!
Found 2 solutions by rfer, faisal rashid:
Answer by rfer(16322)

4/x=11/84+x
11x=336+4x
6x=336
x=56 sea lions
84+x=140 penguins

Answer by faisal rashid(1)

it should be like
4/11 = x/(x+84)
Solving for X gives you 48 sea lions, and penguins are 48+84=132
Checking the answer (48/132)=(4/11)